# Concurrent Composition for Differentially Private Continual Mechanisms

### v2

@@ Appendix A Overview and Modification of [18]
For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $(\epsilon,\delta)$-DP IM w.r.t. a neighboring relation $\sim$. Suppose that $\mathcal{M}$ has discrete answer distributions. Then, for any two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$, the mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ are identical.
Notation A.2.
@@
For short, let $Q=Q_{\mathcal{M}}$ and $A=A_{\mathcal{M}}$. For any $T\in\mathbb{N}$, let $[T]=\{1,\dots,T\}$. For each $t\in\mathbb{N}$ and $b\in\{0,1\}$, we define the function $\mu_{t}^{b}:(Q\times A)^{t}\to[0,1]$ such that $\mu_{t}^{b}(q_{1},\dots,a_{t})$ denotes the probability of $\mathcal{M}(s_{b})$ returning answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$.
@@
Fixing an upper bound $T$ on the number of queries, Lyu [18] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for each $t\in[T]$, satisfying three conditions. They then inductively construct each $\phi_{t}^{b}$ and verify that it satisfies the required conditions. The definition of these conditions is based on a sequence of control function $\mathrm{L}_{t,T}^{0}$ and $\mathrm{L}_{t,T}^{1}$, which have backward recursive definitions for $t$ from $T$ to $1$.
@@
In Section A.1.1, we overview the definition of $\mathrm{L}_{t,T}^{b}$ and its properties. In Section A.1.2, we introduce a new control function $\mathrm{L}_{t}^{b}$ with no dependency on $T$. The goal is to use $\mathrm{L}_{t}^{b}$ instead of $\mathrm{L}_{t,T}^{b}$ in [18]’s proofs. So, in Section A.1.3, we show some properties for $\mathrm{L}_{t}^{b}$, required in later proofs.
In Section A.2, we overview the reduction presented in [18], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$. In this section, we prove the reduction still holds with the new control functions. Finally, in Section A.3, we present a construction for $\phi_{t}$ and show it satisfies all three conditions. We note that this construction is identical in essence to one given in [18], but we extend the proof to work with weaker assumptions.
A.1. Preliminaries
A.1.1. Control Functions $\mathrm{Lower}_{t,T}$ and $\mathrm{L}_{t,T}$.
@@ A.1. Preliminaries
Definition A.3 ($\mathrm{Lower}_{t,T}$).
@@
Claim A.7 ([18]).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$ [^3] In the original claim in [18], $t$ is assumed to be at most $T-1$, which is a typo., and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
@@
Claim A.6 combined with the definition of $\mathrm{L}_{t,T}^{b}$ imply:
Corollary A.8.
@@
As we point out later, in some part of the proof, Lyu [18] wrongly refers to Claim A.7, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
Lemma A.9.
@@
  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\leq e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$.
  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\geq e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then, $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$ as $e^{\epsilon}\geq e^{-\epsilon}$.
@@
where the last inequality holds because for every $q_{t+1}\in Q$, by Claim A.7
@@
A.1.2. Control Function $\mathrm{L}_{t}^{b}$ with is the Limit of $\mathrm{L}_{t,T}$ (Definition).
This section introduces our main new technical contribution. We introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [18]. Thus, differently from [18], we will use $\mathrm{L}_{t}^{b}$ to define the functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$.
However, we first need to argue that the limit of $\mathrm{L}_{t,T}^{b}$ exists. the value of $\mathrm{L}_{t,T}^{b}$ is upper bounded by $1$. We also show that this function is non-decreasing in $T$. As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists. To remove the assumption of having an upper bound $T$ on the number of queries, we introduce the following control function:
Definition A.10 ($\mathrm{L}_{t}^{b}$).
@@
We need the following lemma throughout the section:
Lemma A.11.
@@
Fix $T\geq t$. The proof is by induction on $t$ from $T$ to $1$. For $t=T$, by definition and Lemma A.11, we have
@@
For $t=T$, by definition and Lemma A.11, we have
@@
Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$.
The base case is $t=T$. By Lemma A.11, for every $(q_{1},\dots,a_{t-1},q_{t})$, we have
@@
For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t},q_{t+1})$, we have $\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots,a_{t},q_{t+1})$. We need to show $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$ for every $(q_{1},\dots,a_{t-1},q_{t})$. By definition,
@@
Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$ from $T$ to 00.
The base case is $t=T$. By Lemma A.11, for every $(q_{1},\dots,a_{T})$ we have
@@
For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t+1})$, we have $\mathrm{L}_{t+1,T}^{b}(q_{1},\dots,a_{t+1})\leq\mathrm{L}_{t+1,T+1}^{b}(q_{1},\dots,a_{t+1})$. We need to show $\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t,T+1}^{b}(q_{1},\dots,a_{t})$ for every $(q_{1},\dots,a_{t})$. By definition of $\mathrm{L}_{t,T+1}$ and Lemma A.14, we have
@@
A.1.3. Control Function $\mathrm{L}_{t}^{b}$ (Properties).
In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections.
The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary A.8, Lemma A.9, and Lemma A.13 give the following corollaries, respectively:
Corollary A.16.
@@
Lemma A.19 (Monotone Convergence).
@@
Lemma A.19 combined with Lemma A.15 implies:
Corollary A.20.
For any $b\in\{0,1\}$, $t\in\mathbb{N}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
First, we will show that for any $T>t-1$,
@@
Since $t\neq T$, by definition of $\mathrm{L}_{t-1,T}^{b}$ and Corollary A.5, we have
@@
Thus, by Corollary A.20, we have
@@
To define $\mathcal{P}$, Lyu [18] constructs four families of PMFs $\mathcal{F}_{(0,\bot)}$, $\mathcal{F}_{(1,\bot)}$, $\mathcal{F}_{(0,\top)}$, and $\mathcal{F}_{(1,\top)}$, corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$. Intuitively, each family describes an IM. Specifically, each family consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$, representing the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers.
A.2.1. Constructing $\mathcal{P}$ using $\mathcal{F}_{r}$s.
Roughly speaking, in the interaction between $\mathcal{P}\circ^{*}\mathrm{RR}_{\epsilon,\delta}(b)$ and an adversary $\mathcal{A}$, $\mathcal{P}$ interacts with $\mathrm{RR}_{\epsilon,\delta}(b)$ once and stores its output $r$ as thee first parameter in its state. It then answers the adversary’s queries by itself using the PMFs provided by $\mathcal{F}_{r}$. Specifically, $\mathcal{P}$ has a single initial state $\left(\text{null},()\right)$, where null is an empty field and ()()( ) represents an empty sequence. Let $h$ denote the sequence of interactions between $\mathcal{P}$ and $\mathcal{A}$, which is stored by $\mathcal{P}$ as second parameter of its state. After interacting with $\mathrm{RR}_{\epsilon,\delta}(b)$, $\mathcal{P}$ updates null to $r$. Moreover, upon receiving a query $q$ and before returning an answer $a$, $\mathcal{P}$ appends $q$ and $a$ to the history sequence $h$. To generate a response $a$ to a query $q$, $\mathcal{P}$ samples $a$ from the PMF corresponding to $(h,q)$ in $\mathcal{F}_{r}$, where $h$ is the history of all previous queries and answers.
A.2.2. Constructing $\mathcal{F}_{(b,\bot)}$s.
@@
Lyu [18] constructs a function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in[T]\cup\{0\}$. Let $\phi_{t}^{0}$ and $\phi_{t}^{1}$ denote the functions returning the first and second elements of the output pairs of $\phi_{t}$. For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\bot)}$ is defined based on $\phi_{t}^{b}$s. Intuitively, $\phi_{t}^{b}(q_{1},\dots,a_{t})$ represents the probability that $\mathcal{M}(s_{b})$ returns the answers $a_{1},\dots,a_{t}$ to the fixed queries $q_{1},\dots,q_{t}$. That is, for each query sequence $q_{1},\dots,q_{t}$, $\phi_{t}^{b}(q_{1},\cdot,q_{2},\cdot,\dots,q_{t},\cdot)$ is a PMF over $A^{t}$.
We do not provide the construction of $\phi_{t}$ in this section. However, assuming the existence of $\phi_{t}$ satisfying the following three conditions, we show that for every $b\in\{0,1\}$, the mechanisms $\mathcal{P}\!\circ^{*}\!\mathrm{RR}_{\epsilon,\delta}$ and $\mathcal{M}(s_{b})$ are equivalent. These three conditions were originally stated in [18], with the control function $\mathrm{L}_{t,T}^{b}$ instead of $\mathrm{L}_{t}^{b}$.
For $t=0$, let $\phi_{t}()=(1,1)$. Suppose for every $t\in[T]$, there exists a function $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ satisfying the following conditions for each $b\in\{0,1\}$:
@@
    For any transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
@@
    For any transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
@@
    For any transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\bot)}$ consists of functions $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in\bigcup_{t=1}^{T}(Q\times A)^{t-1}\times Q$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
@@
  • Otherwise, if $\phi_{t-1}(q_{1},\dots,a_{t-1})=0$, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for every answer, except for one fixed arbitrary answer $a^{*}\in A$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$. This case is included only for formality, and the corresponding PMFs are never used by $\mathcal{P}$.
@@
A.2.3. Constructing $\mathcal{F}_{(b,\top)}$s.
Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in[T]\cup\{0\}$. For $t=0$, $\psi_{t}()=(1,1)$. For $t\in[T]$, $\psi_{t}$ is defined as follows:
@@
For each $b\in\{0,1\}$, the family $\mathcal{F}_{(b,\top)}$ consists of functions $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in\bigcup_{t=1}^{T}(Q\times A)^{t-1}\times Q$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
@@
  • Otherwise, $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for every answer, except for one fixed arbitrary answer $a^{*}\in A$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
By condition (II), each $\psi_{t}^{b}$ is non-negative. Also, by Lemma A.11 and condition (III), for every $(q_{1},\dots,a_{t},q_{t})\in(Q\times A)^{t-1}\times Q$,

### v3

@@ 8. Reduction to Interactive Randomized Response
For $\epsilon>0$ and $0\leq\delta\leq 1$, let $\mathcal{M}:S_{\mathcal{M}}\times Q_{\mathcal{M}}\to S_{\mathcal{M}}\times A_{\mathcal{M}}$ be an $(\epsilon,\delta)$-DP IM w.r.t. a neighbor relation $\sim$. Suppose that $\mathcal{M}$ has discrete answer distributions. Then, for every two neighboring initial states $s_{0}$ and $s_{1}$, there exists an IPM $\mathcal{P}$ such that for each $b\in\{0,1\}$:
  • The mechanisms $\mathcal{M}(s_{b})$ and $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ are equivalent.
  • For every $k\in\mathbb{N}$ and every query sequence $(q_{1},\dots,q_{k})\in Q_{\mathcal{M}}^{k}$, if the distributions of answers produced by $\mathcal{M}(s_{0})$ and $\mathcal{M}(s_{1})$ to $(q_{1},\dots,q_{k})$ are identical, then when $\mathcal{P}\!\circ^{*}\!\mathrm{IRR}_{\epsilon,\delta}(b)$ receives the queries $q_{1},\dots,q_{k}$, the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ only once. Otherwise, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}(b)$ at most twice.
Notation 8.1.
@@
For short, let $Q=Q_{\mathcal{M}}$ and $A=A_{\mathcal{M}}$. For each $t\in\mathbb{N}$ and $b\in\{0,1\}$, we define the function $\mu_{t}^{b}:(Q\times A)^{t}\to[0,1]$ such that $\mu_{t}^{b}(q_{1},\dots,a_{t})$ denotes the probability of $\mathcal{M}(s_{b})$ returning answers $a_{1},\dots,a_{t}$ to the queries $q_{1},\dots,q_{t}$.
@@
Fixing an upper bound $T$ on the number of queries, Lyu [19] reduces the construction of $\mathcal{P}$ to defining a sequence of functions $\phi_{t}^{0}$ and $\phi_{t}^{1}$ for $t\in[T]$ that satisfy three conditions. They then inductively construct $\phi_{t}^{b}$ and verify that it satisfies the required conditions. The definition of these conditions is based on a sequence of control function $\mathrm{L}_{t,T}^{0}$ and $\mathrm{L}_{t,T}^{1}$, which have backward recursive definitions for $t$ from $T$ to $1$.
@@
In Section 8.1, we overview the definition of $\mathrm{L}_{t,T}^{b}$ and its properties. In Section 8.1.1, we introduce a new control function $\mathrm{L}_{t}^{b}$ with no dependency on $T$. The goal is to use $\mathrm{L}_{t}^{b}$ instead of $\mathrm{L}_{t,T}^{b}$. In Section 8.1.2, we show some properties for $\mathrm{L}_{t}^{b}$, required in later proofs.
In Section 8.2, we summarize the reduction introduced in [19], with the control functions $\mathrm{L}_{t,T}^{b}$ replaced by $\mathrm{L}_{t}^{b}$. We also modify this reduction by adding a new condition to the original three that the functions $\phi_{t}$ must satisfy. This additional condition is used to limit the number of interactions between $\mathcal{P}$ and $\mathrm{IRR}_{\epsilon,\delta}(b)$. Finally, in Section 8.3, we provide a new construction of $\phi_{t}$ and show that it satisfies all four conditions.
8.1. Control Functions $\mathrm{Lower}_{t,T}$ and $\mathrm{L}_{t,T}$.
@@
Definition 8.2 ($\mathrm{Lower}_{t,T}$).
@@
Claim 8.6 ([19]).
For each $b\in\{0,1\}$, $T\in\mathbb{N}$, $t\in[T]$, [^7] In the original claim in [19], $t$ is assumed to be at most T1T-1, which is a typo. and transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$, we have
@@
Claim 8.5 combined with the definition of $\mathrm{L}_{t,T}^{b}$ imply:
Corollary 8.7.
@@
In some part of the proof, Lyu [19] wrongly refers to Claim 8.6, while the following lemma is required. For completeness, we prove this lemma and use it in the next section.
Lemma 8.8.
@@
  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\leq e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{-\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which equals $B$.
  • If $\mu_{t}^{b}(q_{1},\dots,a_{t})\geq e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t})$, then, $A=\mu_{t}^{b}(q_{1},\dots,a_{t})-e^{\epsilon}\mu_{t}^{1-b}(q_{1},\dots,a_{t}),$ which is at most $B$ as $e^{\epsilon}\geq e^{-\epsilon}$.
@@
where the last inequality holds because for every $q_{t+1}\in Q$, by Claim 8.6
@@
8.1.1. Control Function $\mathrm{L}_{t}^{b}=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}$ (Definition).
This section introduces our main new technical contribution. To remove the assumption of having an upper bound $T$ on the number of queries, we introduce $\mathrm{L}_{t}^{b}$ as limit of $\mathrm{L}_{t,T}^{b}$ and show that it satisfies all properties needed by the proof of [19]. Thus, differently from [19], we will use $\mathrm{L}_{t}^{0}$ and $\mathrm{L}_{t}^{1}$ to define the function $\phi_{t}$.
Definition 8.9 ($\mathrm{L}_{t}^{b}$).
@@
To ensure $\mathrm{L}_{t}^{b}$ is well-defined, we need to show that the limit of $\mathrm{L}_{t,T}^{b}$ exists. We will upper bound the values of $\mathrm{L}_{t,T}^{b}$ by $1$ and show that this function is non-decreasing in $T$. As a result, for each each $t\in\mathbb{N}\cup\{0\}$ and $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$, the limit $\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})$ exists. We need the following lemma throughout the section:
Lemma 8.10.
@@
Fix $T\geq t$. The proof is by induction on $t$ from $T$ to $1$. For $t=T$, by definition and Lemma 8.10, we have
@@
For $t=T$, by definition and Lemma 8.10, we have
@@
Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$: The base case is $t=T$. By Lemma 8.10, for every $(q_{1},\dots,a_{t-1},q_{t})$, we have
@@
For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t},q_{t+1})$, we have
$\mathrm{Lower}_{t+1,T}^{b}(q_{1},\dots,a_{t},q_{t+1})\leq\mathrm{Lower}_{t+1,T+1}^{b}(q_{1},\dots,a_{t},q_{t+1}).$
We need to show $\mathrm{Lower}_{t,T}^{b}(q_{1},\dots,a_{t-1},q_{t})\leq\mathrm{Lower}_{t,T+1}^{b}(q_{1},\dots,a_{t-1},q_{t})$ for every $(q_{1},\dots,a_{t-1},q_{t})$. By definition,
@@
Fix $T\in\mathbb{N}$ and $b\in\{0,1\}$. We show this by reverse induction on $t$ from $T$ to 0.
The base case is $t=T$. By Lemma 8.10, for every $(q_{1},\dots,a_{T})$ we have
@@
For the induction step, assume that $t<T$ and that for every $(q_{1},\dots,a_{t+1})$, we have $\mathrm{L}_{t+1,T}^{b}(q_{1},\dots,a_{t+1})\leq\mathrm{L}_{t+1,T+1}^{b}(q_{1},\dots,a_{t+1})$. We need to show $\mathrm{L}_{t,T}^{b}(q_{1},\dots,a_{t})\leq\mathrm{L}_{t,T+1}^{b}(q_{1},\dots,a_{t})$ for every $(q_{1},\dots,a_{t})$. By definition of $\mathrm{L}_{t,T+1}$ and Lemma 8.13, we have
@@
8.1.2. Control Function $\mathrm{L}_{t}^{b}=\lim_{T\geq t,T\to\infty}\mathrm{L}_{t,T}$ (Properties).
In this section, we present the properties of $\mathrm{L}_{t}^{b}$, required in the next sections. The definition of $\mathrm{L}_{t}^{b}$ combined with Corollary 8.7, Lemma 8.8, and Lemma 8.12 give the following corollaries, respectively:
Corollary 8.15.
@@
Lemma 8.18 (Monotone Convergence).
@@
Lemma 8.18 combined with Lemma 8.14 implies:
Corollary 8.19.
For every $b\in\{0,1\}$, $t\in\mathbb{N}$, and $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
First, we will show that for every T>t1T>t-1,
@@
Since $t-1\neq T$, by definition of $\mathrm{L}_{t-1,T}^{b}$ and Corollary 8.4, we have
@@
Thus, by Corollary 8.19, we have
@@
To define $\mathcal{P}$ and simulate $\mathcal{M}$, Lyu [19] constructs four families of probability mass functions (PMFs), corresponding to each outcome of $\mathrm{RR}_{\epsilon,\delta}$. Intuitively, each family describes an IM. Specifically, each family describes an IM and consists of PMFs over the answer set $A$ for every $t\in[T]$ and every history $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t}\times Q$. These PMFs represent the probability of returning each answer in response to $q_{t}$, conditioned on $q_{1},\dots,a_{t-1}$ being the history of queries and answers. Similarly, we define $\mathcal{P}$ based on four families of PMFs: $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$, corresponding to the four possible output sequences of $\mathrm{IRR}_{\epsilon,\delta}$. We note that in our construction, each family includes a PMF over $A$ for every $t\in\mathbb{N}$ and every history $(q_{1},\dots,a_{t-1},q_{t})$. Moreover, the definitions of these PMFs are not the same as the ones in [19].
8.2.1. Constructing $\mathcal{P}$.
In [19], the post-processing mechanism initially interacts with $\mathrm{RR}_{\epsilon,\delta}$ and stores its output $r=(\tau,c)\in\{\top,\bot\}\times\{0,1\}$. It then answers the adversary’s queries by itself using the family of PMFs corresponding to $r$. Our high-level idea is to construct the families $\mathcal{F}_{(\bot,0)}$, $\mathcal{F}_{(\bot,1)}$, $\mathcal{F}_{(\top,0)}$, and $\mathcal{F}_{(\top,1)}$ in such a way that, under certain conditions, knowledge of $\tau$ alone suffices to choose the appropriate PMF. Recall that the IPM $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ instead of $\mathrm{RR}_{\epsilon,\delta}$. We design $\mathcal{P}$ to initially interact with $\mathrm{IRR}_{\epsilon,\delta}$ and store its response. $\mathcal{P}$ then interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time when it cannot choose a PMF using only the first response.
[…]
To generate a response $a$ to query $q$, $\mathcal{P}$ proceeds as follows: If $q$ is the first query (i.e., $r=()$), then $\mathcal{P}$ initially interacts with $\mathrm{IRR}_{\epsilon,\delta}$ and appends its response to $r$. After this, whether $q$ is the first query or not, when the following condition does not hold for the first time, $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$ for the second time and appends its answer to $r$: the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\top,0)}$ and $\mathcal{F}_{(\top,1)}$ are identical, and the PMFs corresponding to $(h,q)$ in $\mathcal{F}_{(\bot,0)}$ and $\mathcal{F}_{(\bot,1)}$ are identical. After that, $\mathcal{P}$ never interacts with $\mathrm{IRR}_{\epsilon,\delta}$ again.
After the potential interactions with $\mathrm{IRR}_{\epsilon,\delta}$, $\mathcal{P}$ generates the answer $a$ as follows:
  • If $r\in\{\top,\bot\}\times\{0,1\}$, then $\mathcal{P}$ samples $a$ from the PMF corresponding to $(h,q)$ in $\mathcal{F}_{r}$.
  • Otherwise, by design, $r\in\{\top,\bot\}$, and the PMFs in $\mathcal{F}_{(r,0)}$ and $\mathcal{F}_{(r,1)}$ corresponding to $(h,q)$ are identical. In this case, $\mathcal{P}$ samples $a$ from this common PMF.
8.2.2. Constructing $\mathcal{F}_{(\bot,b)}$s.
Using the control function $\mathrm{Lower}_{t,T}$, Lyu [19] introduces three conditions and constructs a sequence of functions $\phi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for $t\in\{1,\dots,T\}$ that satisfy these conditions. The families of PMFs are defined then based on these functions. We remove the dependence on the bound $T$ by replacing $\mathrm{Lower}_{t,T}$ with $\mathrm{L}_{t}$ in the conditions, and extend the construction of $\phi_{t}$ to all $t\in\mathbb{N}$. Furthermore, we introduce an additional condition to those given in [19], which we will later use to bound the number of times $\mathcal{P}$ interacts with $\mathrm{IRR}_{\epsilon,\delta}$.
@@
Let $\phi_{t}^{0}$ and $\phi_{t}^{1}$ denote the functions that return the first and second components, respectively, of the output of $\phi_{t}$. Intuitively, each $\phi_{t}^{b}$ represents an IM: for every query sequence $q_{1},\dots,q_{t}$, it assigns a probability to each answer sequence $a_{1},\dots,a_{t}$. That is, for any query sequence $q_{1},\dots,q_{t}$, the function $\phi_{t}^{b}(q_{1},\cdot,q_{2},\cdot,\dots,q_{t},\cdot)$ defines a PMF over $A^{t}$.
For $t=0$, we define $\phi_{t}()=(1,1)$. We defer the detailed construction of the functions $\phi_{t}$ to Section 8.3. However, we assume the existence of $\phi_{t}$ satisfying the following conditions for each $b\in\{0,1\}$:
@@
    For every transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
@@
    For every transcript $(q_{1},\dots,a_{t})\in(Q\times A)^{t}$,
@@
    For every transcript $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{t-1}\times Q$,
@@
  4. (IV)
    For each query sequence $(q_{1},\dots,q_{t})\in Q^{t}$, if $\mu_{t}^{0}(q_{1},\dots,a_{t})=\mu_{t}^{1}(q_{1},\dots,a_{t})$ for every answer sequence $(a_{1},\dots,a_{t})\in A^{t}$, then
    $\phi_{t}^{0}(q_{1},\dots,a_{t})=\phi_{t}^{1}(q_{1},\dots,a_{t})\geq\min\{\mu_{t}^{0}(q_{1},\dots,a_{t}),\mathrm{L}_{t}^{0}(q_{1},\dots,a_{t})+\mathrm{L}_{t}^{1}(q_{1},\dots,a_{t})\}$
    for every $(a_{1},\dots,a_{t})\in A^{t}$.
For each $b\in\{0,1\}$, the family $\mathcal{F}_{(\bot,b)}$ consists of functions $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ defined for every $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{*}\times Q$, where $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
@@
  • Otherwise, if $\phi_{t-1}(q_{1},\dots,a_{t-1})=0$, $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for all $a_{t}\in A$, except for one fixed arbitrary answer $a^{*}\in A$, for which $f_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$. This case is included only for formality, as the corresponding PMFs are never used by $\mathcal{P}$.
@@
8.2.3. Constructing $\mathcal{F}_{(\top,b)}$s.
Similar to the previous section, $\mathcal{F}_{(0,\top)}$ and $\mathcal{F}_{(1,\top)}$ are defined based on a sequence of functions $\psi_{t}:(Q\times A)^{t}\to[0,1]^{2}$ for each $t\in\mathbb{N}\cup\{0\}$. For $t=0$, $\psi_{t}()=(1,1)$. For $t\in\mathbb{N}$, we define $\psi_{t}$ as follows:
@@
For each $b\in\{0,1\}$, the family $\mathcal{F}_{(\top,b)}$ consists of functions $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}:A\to[0,1]$ for each $(q_{1},\dots,a_{t-1},q_{t})\in(Q\times A)^{*}\times Q$, where $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}$ is defined as follows:
@@
  • Otherwise, $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a_{t})=0$ for all $a_{t}\in A$, except for one fixed arbitrary answer $a^{*}\in A$, for which $g_{q_{1},\dots,a_{t-1},q_{t}}^{b}(a^{*})=1$.
By condition (II), $\psi_{t}^{b}$ is non-negative. Also, by Lemma 8.10 and condition (III),